One number is 2 times a first number. A third number is 100 more than the first number. If the sum of the three numbers is 300​, find the numbers.
Answer by fractalier(6550)

Call the first number x. The second number is 2x. The third number is x+100. If they add to 300 we write
x + 2x + x+100 = 300
4x + 100 = 300
4x = 200
x = 50
2x = 100
x+100 = 150